OpenStudy (priyar):

if the general solution of the diff. eq. y'=y/x +f(x/y), for some function y,is given by yln|cx|=x ,where c is an arbitary constant,then f(2)=?

hartnn (hartnn):

tried this? find y' from ln|cx|=x plug it in diff. eq. put x = 2y

OpenStudy (priyar):

the variable y remains in the eq..

hartnn (hartnn):

y' ln |cx| + y/x = 1 y' = (1-y/x)/ ln |cx| but ln |cx| = y/x y' = (1-y/x)/(y/x) y' = y/x + f(x/y) (1-y/x)/(y/x) = y/x + f(x/y) x= 2y, y/x = 1/2 (1-1/2)/(1/2) = 1/2 + f(2)

OpenStudy (priyar):

oh ok i didn't substitute..ln|cx|=y/x..

OpenStudy (priyar):

thank u @hartnn !

hartnn (hartnn):

welcome ^_^